Consider the reaction:

H2(g) + I2(s)    =    2HI (g)

Given an initial mass of 19.02g H2 an excess of I2 and assuming that all of the reactant is conserved to products and non is lost calculate the mass (g) of HI produced by the reaction?
Answer #1

Sol .

Reaction : H2(g) + I2(g) ---> 2HI(g)

As mass of H2 = 19.02 g

Molar mass of H2 = 2.02 g/mol

So , moles of H2 = mass of H2 / molar mass of H2 = 19.02 / 2.02 = 9.4158 mol

Now , from reaction , 1 mole of H2 produces 2 moles of HI .

So , 9.4158 moles of H2 produces = 9.4158*2 = 18.8316 moles of HI

As molar mass of HI = 127.91 g / mol

Therefore , mass of HI produced = Moles of HI produced * Molar mass of HI

= 18.8316*127.91

= 2408.75 g